There is some non-trivial sign flipping in the sensors/coils in this new setup because it is a hybrid one with the old interfacing electronics (D000210, D010001) and the new Satellite Amplifier (D080276). So I haven't yet gotten the damping working. I am leaving the PSL shutter closed and will keep working on this today/tomorrow. I have made various changes to the c1mcs realtime model and the c1susaux database record where MC1 is concerned. I have backups of the old ones so we can always go back to that if we so desire.
In the meantime, the PSL shutter is closed and there is no light to the IFO.
Update 1700: I've implemented some basic damping and now the IMC is now locked. The WFS loop runs away when I enable it, probably some kind of weird interaction with the (as of now untuned) MC1 local damping loops. I will write up a more detailed report later.
Update 2300: Did the following:
Dropping this for tonight, I'll continue tomorrow. Meanwhile, the OSEM input matrix measurement is being repeated overnight. PSL shutter is closed.